OpenStudy (anonymous):

Maple Grove wants to include two types of maple trees, the namesake of the city,for their parks.Two varieties of maple trees have been selected.One variety costs $80 per tree; the other more colorful variety costs $85 per tree. The tree budget must not exceed $75,000. Citizens want more of the colorful $85 trees than the others if possible. Let x be the number of $80 trees and y the number of $85 trees. Which system of inequalities represents the situation?

OpenStudy (saifoo.khan):

LANA!

OpenStudy (anonymous):

this question is a waste of time!!!!!!!

OpenStudy (saifoo.khan):

prolly.

OpenStudy (lalaly):

SAIF:D

OpenStudy (saifoo.khan):

LARA!

OpenStudy (lalaly):

SAIFOO

OpenStudy (saifoo.khan):

LALA.

OpenStudy (anonymous):

can i get some help?

OpenStudy (anonymous):

lala!!!!! lolx

OpenStudy (lalaly):

Sword! loool

OpenStudy (saifoo.khan):

LALAY!

OpenStudy (anonymous):

where are the options

OpenStudy (anonymous):

???????????????????????????

OpenStudy (saifoo.khan):

are u cesar or nadia?

OpenStudy (lalaly):

hahahaha

OpenStudy (saifoo.khan):

LOL.

OpenStudy (anonymous):

are u a julio cesar

OpenStudy (anonymous):

????????????????????

OpenStudy (anonymous):

where is desdemona

OpenStudy (anonymous):

shakespear has written a good play on u

OpenStudy (saifoo.khan):

http://en.wikipedia.org/wiki/Julio_Iglesias

OpenStudy (anonymous):

80x+85y>= 75,000 x<=y 80x+85y>=75,000 x>=y 80x+85y<=75,000 x>=y 80x+85y<=75,000 x<=y

OpenStudy (anonymous):

those are the options